### v1

@@ Abstract
With the rise of data volume and computing power, seismological research requires more advanced skills in data processing, numerical methods, and parallel computing. We present the experience of conducting training workshops over various forms of delivery to support the adoption of large-scale High-Performance Computing and Cloud computing to advance seismological research. The seismological foci were on earthquake source parameter estimation in catalogs, forward and adjoint wavefield simulations in 2 and 3 dimensions at local, regional, and global scales, earthquake dynamics, ambient noise seismology, and machine learning. This contribution describes the series of workshops, the learning outcomes of the participants, and lessons learned by the instructors. Our curriculum was grounded on open and reproducible science, large-scale scientific computing and data mining, and computing infrastructure (access and usage) for HPC and the cloud. We also describe the types of teaching materials that have proven beneficial to the instruction and the sustainability of the program. We propose guidelines to deliver future workshops on these topics.
@@ 1 Introduction
Numerical methods and the rise of high-performance computing have fueled the development of computational seismology, notably to solve the wave equation in complex media (e.g., Komatitsch and Vilotte, 1998; Komatitsch et al., 2002) and coupled to complex source models for purposes of physics-based ground-motion simulations (e.g., Graves et al., 2011) and for seismic imaging (e.g., Liu and Gu, 2012; Tromp, 2020). As examples, the SPECFEM package (Komatitsch and Tromp, 2002a; Komatitsch et al., 2004) have supported a new era of passive-source (earthquake, ambient noise) full-waveform-inversion (FWI) (e.g., Tape et al., 2009; Peter et al., 2011; Bozdağ et al., 2016; Chow et al., 2020) and the SeisSol software enables realistic simulations of 3D earthquake rupture dynamics (e.g., Käser et al., 2010; Pelties et al., 2012, 2014; Krenz et al., 2021; Gabriel et al., 2023; Uphoff et al., 2024).
Big data seismology is also vastly expanding, as continuous seismic data is recorded by hundreds or thousands of seismic stations worldwide. New methods emerged to include array processing (e.g., Rost and Thomas, 2009), ambient field (noise) seismology (e.g., Nakata et al., 2019), and machine learning (e.g., Kong et al., 2019; Mousavi and Beroza, 2022). Discoveries of new tectonic and environmental phenomena invigorate the collection of large seismic data sets, leading to an exponential growth in data volumes and bringing our community to an era of petabyte-scale archives (Arrowsmith et al., 2022). Novel computing infrastructures such as cloud computing are particularly well suited for big data seismological research (MacCarthy et al., 2020; Krauss et al., 2023; Ni et al., 2023).
The broad adoption of open-source software based on Python (e.g., Beyreuther et al., 2010) or Julia (e.g., Jones et al., 2020), as well as version control hosted on GitHub, Bitbucket, and GitLab is transforming research practice and standards (Chue Hong et al., 2022; Barker et al., 2022). Scientific journals require publicly hosted repositories or software availability, and Jupyter notebooks have become a popular avenue for sharing simple workflows (Pérez and Granger, 2007; Pimentel et al., 2019).
@@
The COVID-19 pandemic has transformed education: students and teachers had to transition from in-person to remote, online learning. Several efforts have contributed to improving remote access to seismology education, such as the ROSES program (Brudzinski et al., 2021). These contributions have democratized education through pedagogical approaches analyzing small datasets, using approximate solutions, or performing modest simulations using single nodes and Python notebooks. However, a gap remains in the adoption of advanced computing platforms, such as high-performance computing (HPC) infrastructure and cloud computing. This article presents recent developments by the project SCOPED (Seismological COmputational Platform to Empower Discovery Tape et al. (2022); Wang et al. (2023); Denolle et al. (2024)) and collaborations with other projects (e.g., the Statewide California Earthquake Center (SCEC), EarthScope, and the European projects Geo-Inquire, DT-GEO and ChEESE-2P (Folch et al., 2023)) to help close that gap for students and researchers with multi-modal educational efforts.
The goal of the SCOPED project is to develop a cyber-infrastructure that enables hybrid model–data research in seismology by utilizing cloud and HPC infrastructures, open-source software, and containerization. Research enabled by SCOPED includes 1) AI-enhanced earthquake source characterization and catalog building, 2) full-waveform inversion for source mechanisms, 3) full-waveform inversion for Earth imaging across scales, and 4) time-lapse imaging of the subsurface. The SCOPED community expressed their research interest, which we illustrate with Fig. 1. This article discusses the workshops held as part of the SCOPED project (Table 1), and in particular by its use of containers.
@@ 2 A Broad Survey of the Seismology Community
We ran multiple surveys to gauge the community interest in using widely employed seismological software in advanced computing environments such as HPC and the Cloud in anticipation of workshops between spring 2022 and spring 2024. The workshops were delivered in virtual, hybrid, and in-person formats. Surveys and workshops were announced in domain-specific mailing lists, such as the EarthScope Consortium and SCEC, and social media platforms (X formerly Twitter and LinkedIn). The surveys were tailored to each workshop, and the data collected mainly focused on familiarity with the required technical skills. Overall, we received 976 responses, although some may come from the same individuals. Our workshops had a total of 574 participants, with over 130 joining in-person. Some of the surveys presented here has a broad community reach, and our post-event surveys only had the workshop participants. We found that the timing and frequency of surveys had an effect on the response rate. Post-event surveys were successful only if participants completed them during the event. Due to differences in response rates for pre- and post-event surveys, our analysis combines common questions and categories from both types. The response rate was above 96% for requests during the workshop, whether the meeting was in person or virtual, while it was 13% in the case of the 2024 SSA workshop.
Survey questions were designed to minimize the imposter syndrome as suggested by Huppenkothen et al. (2018). For instance, we asked participants about their familiarity with shell scripting in various forms: “How familiar are you with computing programming from a command line (i.e., within a terminal window)?" with the response fields of “No experience, Some Experience, Extensive Experience." We also asked their familiarity with version control was assessed with questions such as “All of my active research projects over the past year are on GitHub with many check-ins". Another example to assess their proficiency in Python was “I use Python in my life" with the multiple-choice answer “several hours a week and mostly in the classroom", “several hours a day in my research", “all and every day!", and “Never-ever". We also gathered preliminary knowledge about the technical skill levels of the survey respondents. We emphasize that our surveys were a “self-assessment," which likely provided a biased response. To limit these effects, we employed strategies similar to Huppenkothen et al. (2018). For example, we assessed version control (e.g., GitHub) familiarity by asking how many research project participants were hosted on the platform. Another example is assessing Python familiarity by asking how many hours a week a participant uses Python.
@@
Figure 2: Proportion of participants as a summary over 7 workshop surveys for a) career levels and b) self-assessed proficiency in computing tools (CLI=command line interface).
We use three skill-based proficiency levels as a relative metric for the seismological community in Figure 2b. Overall, most participants felt comfortable with CLI tools. 90% of the participants reported having a sufficient level of familiarity with Jupyter notebooks and Docker. 40% of the participants declared being sufficiently experts in Python for their research, though a majority declared having a medium level of familiarity. Interestingly, version control using Git ranks last in our assessment, as most participants report having a medium level of comfort, and 16% declare having no experience with GitHub.
@@ 3.1 CI on HPC
The main computing environment benefiting seismology today is HPC, which is enabled by clusters of thousands of tightly connected nodes managed by large computing centers, such as the Texas Advanced Computing Center (TACC), the San Diego Supercomputing Center (SDSC). Clusters are designed for parallelized workflows that require large shared memory and up to thousands of nodes. They require optimized software with demonstrated efficient performance scaling and proficiency in job scheduler-memory management and storage architecture. HPC centers often provide training.
Over the course of the workshops, we have trained participants in various aspects of HPC. The lectures entailed training on the fundamentals of HPC, how to write allocation proposals for HPC resources, and the parallelization of workflows leveraging shared or distributed memory architectures. We also trained a few selected groups of participants to access and run forward and adjoint simulations to compute 3D synthetic seismograms and data sensitivity kernels with SPECFEM3D_GLOBE (Komatitsch and Tromp, 2002a, b) for FWI and dynamic rupture simulations with SeiSol (Käser et al., 2010) on the Frontera system (Stanzione et al., 2020) at the Texas Advanced Computing Center (TACC).
@@ 3.2 CI on Cloud Computing
We trained participants in cloud computing concepts, such as its design to interact with storage and perform large-scale deployments, which is still in its infancy in seismology, and user access remains a challenge (Krauss et al., 2023). We presented diverse strategies for using cloud resources. They accessed Google Colab notebooks provided by Google Cloud Platform, which are pre-configured Python-based jupyterhubs, and learned how to customize them by manually installing additional dependencies. Accessibility is a major benefit of the Colab approach, as VM specifications can easily be modified on the Google Colab web interface. The free version is limited in size.
The SCOPED project chose AWS as the cloud provider due to the availability of large seismic datasets already hosted on AWS Simple Storage Service (S3). The workshop covered 1) various ways to access AWS cloud resources, 2) how to launch an AWS computing resource on the Elastic-Computing (EC2) referred to as an instance from scratch via the web console, 3) how to install basic research software into their instances, and 4) how to run research-grade problems on the Cloud. We used the typical AWS web console to deploy compute resources during one of our workshops and illustrated it in Figure 3. We taught popular tools for software environments and versioning, such as git , Docker, and conda. It should be noted that significant effort was required to simplify and prepare instructions for streamlined access and use of AWS instances. In particular, it is not trivial to open and access a jupyter lab, and we curated the training materials to achieve this (HPS).
Additionally, we taught various ways to conduct research workflows on the cloud: cloud-native workflows that incorporate cloud services as part of the design (e.g., NoisePy,Jiang and Denolle (2020), HPS), and, alternatively, workflows that are lifted to the cloud, e.g the Lamont-Doherty Earth Observatory earthquake catalog production workflow (Wang et al., 2024) that includes algorithms for event detection and phase arrival time measurements (QuakeFlow , Zhu et al. (2023)), discrimination (SpecUFEx, Holtzman et al. (2018); Sawi et al. (2022) ), and relocation (HypoDD ,Waldhauser and Ellsworth (2000)).
@@ 3.3 Open-Source and Containerized Software
The SCOPED platform gathers open-source software that tackles big data and large-scale simulation research. When writing this paper, SCOPED includes full waveform modeling and inversion, AI-aided earthquake catalog building and source characterization, ambient field seismology, and earthquake dynamic rupture simulations.
The underpinning strategy for deploying our software is containerization, which enhances portability and exploits negligible computing overhead (Wang et al., 2019) once successfully containerized. Containers are isolated images of software and its dependencies that can be deployed on various operating systems and hardware (Docker, Singularity). To grow our user and developer community, SCOPED flagship software is containerized with tutorials provided in the form of Jupyter notebooks. We developed a SeisSCOPED container registry in which the container base holds minimum dependencies. Additional dependencies can be added to the container base: for instance, an HPC-specific container loads modules for Message-Processing Interface - MPI, and a cloud-specific container has cloud-provider Command-Line-Interface CLI-specific packages. One major advantage of the containerized software approach is long-term stability; workshop users and future students alike are able to leverage the same workshop container and its pinned software dependencies, training materials, and test data. Another powerful use of notebooks is integrating shell scripting within notebook cells using system commands. As an example, one can deploy parallelized Python scripts on Azure Pool (Krauss et al., 2023), AWS Batch resources in a single notebook, or run parallelized SPECFEM simulations through JupyterLab (HPS).
While high-performance computing favors the use of scripting, compiled executables, and minimal container sizes, training materials benefit from attaching small test data and notebooks for documentation and visualization of results. Containers may add IPython and ipykernel dependencies to support Jupyter notebooks and small test data to a given container. Opening Jupyter notebooks from remote servers can pose a challenge in group settings, especially on cloud instances. Throughout workshops, our team scripted the following command line to easily allow access to Jupyter notebooks from a container by fixing the token and IP address:
@@
where IdentityProvider.token=scoped gives a specific token (this avoids users tracking it in the long logs printed on the terminal), allow-root grants root access for users inside the container volume, ip=0.0.0.0 tells the server to listen on all available network interfaces, and nohup jupyter lab --no-browser & opens JupyterLab without launching the browser, which is not possible on remote virtual machines protected by a private IP address. This small code snippet was designed to accelerate research rather than being hung up on infrastructure.

### v2

@@ Abstract
With the rise of data volume and computing power, seismological research requires more advanced skills in data processing, numerical methods, and parallel computing. We present the experience of conducting training workshops over various forms of delivery to support the adoption of large-scale High-Performance Computing and Cloud computing to advance seismological research. The seismological foci were on earthquake source parameter estimation in catalogs, forward and adjoint wavefield simulations in 2 and 3 dimensions at local, regional, and global scales, earthquake dynamics, ambient noise seismology, and machine learning. This contribution describes the series of workshops that were delivered as part of research projects, the learning outcomes of the participants, and lessons learned by the instructors. Our curriculum was grounded on open and reproducible science, large-scale scientific computing and data mining, and computing infrastructure (access and usage) for HPC and the cloud. We also describe the types of teaching materials that have proven beneficial to the instruction and the sustainability of the program. We propose guidelines to deliver future workshops on these topics.
@@ 1 Introduction
Numerical methods and the rise of high-performance computing have fueled the development of computational seismology, notably to solve the wave equation in complex media (e.g., Komatitsch and Vilotte, 1998; Komatitsch et al., 2002; Bao et al., 1998; Olsen and Archuleta, 1996; Graves, 1998) and coupled to complex source models for purposes of physics-based ground-motion simulations (e.g., Graves et al., 2011) and for seismic imaging (e.g., Liu and Gu, 2012; Tromp, 2020). As examples, the open-source SPECFEM package (e.g., Komatitsch and Tromp (2002a); Komatitsch et al. (2004), https://specfem.org/) has supported a new era of passive-source (earthquake, ambient noise) full-waveform-inversion (FWI) (e.g., Tape et al., 2009; Peter et al., 2011; Bozdağ et al., 2016; Chow et al., 2020) and the open-source SeisSol software (https://seissol.org/) enables realistic simulations of 3D earthquake rupture dynamics (e.g., Käser et al., 2010; Pelties et al., 2012, 2014; Krenz et al., 2021; Gabriel et al., 2023; Uphoff et al., 2024).
Big data seismology is also vastly expanding, as continuous seismic data is recorded by more and more permanent stations worldwide, tens of thousands at the time of writing, and many more in temporary deployments. New methods emerged to include array processing (e.g., Rost and Thomas, 2009), ambient field (noise) seismology (e.g., Nakata et al., 2019), and machine learning (e.g., Kong et al., 2019; Mousavi and Beroza, 2022). Discoveries of new tectonic and environmental phenomena invigorate the collection of large seismic data sets, leading to an exponential growth in data volumes and bringing our community to an era of petabyte-scale archives (Arrowsmith et al., 2022). Novel computing infrastructures such as cloud computing are particularly well suited for big data seismological research (MacCarthy et al., 2020; Krauss et al., 2023; Ni et al., 2023).
The broad adoption of open-source software in seismology based on Python (e.g., Beyreuther et al., 2010) or Julia (e.g., Jones et al., 2020), as well as version control hosted on GitHub, Bitbucket, and GitLab is transforming research practice and standards (Chue Hong et al., 2022; Barker et al., 2022). Scientific journals require publicly hosted repositories or software availability. The Jupyter project (Pérez and Granger, 2007; Pimentel et al., 2019) encompasses a suite of interactive computing tools: JupyterLab, a modern, integrated development environment that unifies notebooks, code editors, and more; Jupyter Notebook, the classic, document-focused interface for interactive computing; and JupyterHub, a server that enables multi-user access to these environments.
@@
The COVID-19 pandemic has transformed education: students and teachers had to transition from in-person to remote, online learning. Several efforts have contributed to improving remote access to seismology education, such as the ROSES program (Brudzinski et al., 2021). These contributions have democratized education through pedagogical approaches analyzing small datasets, using approximate solutions, or performing modest simulations using single nodes and Jupyter Notebooks. However, a gap remains in the adoption of advanced computing platforms, such as high-performance computing (HPC) infrastructure and cloud computing. This article presents recent developments by the project SCOPED (Seismological COmputational Platform to Empower Discovery Tape et al. (2022); Wang et al. (2023); Denolle et al. (2024)) and collaborations with other projects (e.g., the Statewide California Earthquake Center (SCEC), EarthScope, and the European projects Geo-Inquire, DT-GEO and ChEESE-2P (Folch et al., 2023)) to help close that gap for students and researchers with multi-modal educational efforts.
The goal of the SCOPED project is to develop a cyber-infrastructure that enables hybrid model–data research in seismology by utilizing cloud and HPC infrastructures, open-source software, and containerization. Software containerization is a lightweight virtualization of software and its dependencies into a portable, isolated environment, ensuring consistency across different computing environments. Research enabled by SCOPED includes 1) machine-learning-enhanced earthquake source characterization and catalog building, 2) full-waveform inversion for source mechanisms, 3) full-waveform inversion for Earth imaging across scales, and 4) time-lapse imaging of the subsurface. The SCOPED community expressed their research interest, which we illustrate with Fig. 1. This article discusses the workshops held as part of the SCOPED project (Table 1), and in particular, by its use of containers. The main goal was for workshop participants to learn about research software and how to access and use high-performance computing resources, clusters from HPC centers and resources from Cloud.
@@ 2 A Broad Survey of the Seismology Community
As part of SCOPED, we ran multiple surveys to gauge the community interest in using widely employed seismological software in advanced computing environments such as HPC and the Cloud before the workshops (see Table 1.) The workshops were delivered in virtual, hybrid, and in-person formats with a target audience of researchers with a skill level of advanced graduate levels in seismology. Surveys and workshops were announced in domain-specific mailing lists, such as the EarthScope Consortium and SCEC, and social media platforms (X, formerly Twitter and LinkedIn). The surveys were tailored to each workshop, and the data collected mainly focused on familiarity with the required technical skills. Overall, we received 976 responses, although some may come from the same individuals. Our workshops had a total of 574 participants, with over 130 joining in person. Some of the surveys presented here have a broad community reach, and our post-event surveys only had the workshop participants. We found that the timing and frequency of surveys had an effect on the response rate. Post-event surveys were successful only if participants completed them during the event. Due to differences in response rates for pre- and post-event surveys, our analysis combines common questions and categories from both types. The response rate was above 96% for requests during the workshop, whether the meeting was in person or virtual, while it was 13% in the case of the 2024 SSA workshop.
Survey questions were designed to minimize the imposter syndrome as suggested by Huppenkothen et al. (2018). For instance, we asked participants about their familiarity with shell scripting in various forms: “How familiar are you with computing programming from a command line (i.e., within a terminal window)?" with the response fields of “No experience, Some Experience, Extensive Experience." We also asked about their familiarity with version control with questions such as “All of my active research projects over the past year are on GitHub with many check-ins". Another example to assess their proficiency in Python was “I use Python in my life" with the multiple-choice answer “several hours a week and mostly in the classroom", “several hours a day in my research", “all and every day!", and “Never-ever". We also gathered preliminary knowledge about the technical skill levels of the survey respondents. We emphasize that our surveys were a “self-assessment," which likely provided a biased response.
@@
Figure 2: Proportion in percent (%) of participants as a summary over 7 workshop surveys for a) career levels and b) self-assessed proficiency in computing tools (CLI=command line interface).
We use three skill-based proficiency levels as a relative metric for the seismological community readiness for working with our large-scale software in Figure 2b. Overall, most participants felt comfortable with command-line interface (CLI) tools. 90% of the participants reported having a sufficient level of familiarity with running Jupyter Notebooks and using Docker images. 40% of the participants declared being sufficiently experts in Python for their research, though a majority declared having a medium level of familiarity. Interestingly, version control using Git ranks last in our assessment, as most participants report having a medium level of comfort, and 16% declare having no experience with GitHub.
@@ 3.1 CI on HPC
The primary computing environment benefiting large-scale seismology today is HPC, which is enabled by clusters of thousands of tightly connected nodes managed by large computing centers, such as the resources used in our workshops, the Texas Advanced Computing Center (TACC) and the San Diego Supercomputing Center (SDSC). Clusters are designed for parallelized workflows. Many seismological applications, such as full-waveform inversion, seismic imaging, and earthquake cycle simulations, follow a Single-Program Multiple Data model, where the same code runs across multiple nodes, each processing a different subset of the data. These HPC workflows require optimized software with efficient parallel scaling, as well as proficiency in job scheduling, memory management, and storage architecture. HPC centers often provide training to help researchers develop and optimize these computational techniques.
Over the course of the workshops, we have trained participants in various aspects of HPC. The lectures entailed training on the fundamentals of HPC, how to write allocation proposals for HPC resources (e.g., what are the elements to show in a proposal to an NSF access proposal, or how to get AWS education or cloudbank), and the parallelization of workflows leveraging shared or distributed memory architectures. We also trained a few selected groups of participants, approximately 80 total, to access and run forward and adjoint simulations to compute 3D synthetic seismograms and data sensitivity kernels with SPECFEM3D_GLOBE (Komatitsch and Tromp, 2002a, b) for FWI and dynamic rupture simulations with SeiSol (Käser et al., 2010) on the Frontera system (Stanzione et al., 2020) at the Texas Advanced Computing Center (TACC).
@@ 3.2 CI on Cloud Computing
Cloud computing is still in its infancy in seismology, and user access remains a challenge (Krauss et al., 2023). We trained participants in cloud computing concepts, such as its design to interact with storage and perform large-scale deployments. We presented diverse strategies for using cloud resources to workshop participants. They accessed Google Colab (https://colab.research.google.com/) provided by Google Cloud Platform, which is a pre-configured Python-based Jupyter Notebook, and learned how to customize them by manually installing additional dependencies. Accessibility is a major benefit of the Colab approach, as VM specifications can easily be modified on the Google Colab web interface. The free version of Google Colab is limited in size (e.g., a few CPUs, 12 GB of RAM, and 50 GBs of storage).
The SCOPED project chose AWS as the cloud provider due to the availability of large seismic datasets already hosted on AWS Simple Storage Service (S3) (Northern and Southern California data centers as NCEDC and SCEDC NCEDC (2014); Yu et al. (2021)). The workshop covered 1) various ways to access AWS cloud resources, 2) how to launch an AWS computing resource on the Elastic-Computing (EC2) referred to as an instance from scratch via the web console, 3) how to install basic research software into their instances, and 4) how to run research-grade problems on the Cloud. We used the typical AWS web console to deploy compute resources during one of our workshops and illustrated it in Figure 3. We taught basic concepts and practiced popular tools for software environments and versioning, such as git, Docker, and conda. We note that significant effort was required to simplify and prepare instructions for streamlined access and use of AWS instances. In particular, it is not trivial to open and access a Jupyter Lab, and we curated the training materials to achieve this in our Jupyter Book (HPS; High Performance Seismology (SeisSCOPED, 2024)).
Additionally, we taught various ways to conduct research workflows on the cloud: cloud-native workflows that incorporate cloud services as part of the design (e.g., NoisePy, Jiang and Denolle (2020), HPS), and, alternatively, workflows that are lifted-and-shifted migrated to the cloud, e.g the Lamont-Doherty Earth Observatory earthquake catalog production workflow (Wang24) that includes algorithms for event detection and phase arrival time measurements (QuakeFlow, Zhu et al. (2023)), discrimination (SpecUFEx, Holtzman et al. (2018); Sawi et al. (2022) ), and relocation (HypoDD, Waldhauser and Ellsworth (2000)).
@@ 3.3 Open-Source and Containerized Software
The SCOPED platform gathers open-source software that tackles big data and large-scale software research. Currently, SCOPED includes full waveform modeling and inversion, machine-learning-aided earthquake catalog building and source characterization, ambient field seismology, and earthquake dynamic rupture simulations.
The underpinning strategy for deploying our software is containerization, which enhances portability and exploits negligible computing overhead (Wang et al., 2019) once successfully containerized. Containers are isolated images of software and its dependencies that can be deployed on various operating systems and hardware (Docker, Singularity). Containers promote long-term sustainability and reproducibility of the computing analysis. To grow our user and developer community, SCOPED flagship software is containerized with tutorials provided in the form of Jupyter Notebooks. We developed a SeisSCOPED container registry in which the container base holds minimum dependencies. Additional dependencies can be added to the container base: for instance, an HPC-specific container loads modules for Message-Processing Interface - MPI, and a cloud-specific container has cloud-provider Command-Line-Interface CLI-specific packages. One significant advantage of the containerized software approach is long-term stability; workshop users and future students alike can leverage the same workshop container and its pinned software dependencies, training materials, and test data. Another powerful use of notebooks is integrating shell scripting within notebook cells using system commands. As an example, one can deploy parallelized Python scripts on Azure Pool (Krauss et al., 2023), AWS Batch resources in a single notebook, or run parallelized SPECFEM simulations through Jupyter Lab (HPS).
While high-performance computing favors the use of scripting, compiled executables, and minimal container sizes, training materials benefit from attaching small test data and notebooks for documentation and visualization of results. Containers may add IPython and ipykernel dependencies to support Jupyter Notebooks and small test data to a given container. Especially for cloud computing, opening Jupyter Notebooks from remote servers can pose a challenge in group settings, especially on cloud instances. Throughout workshops, our team came up with the following command line to easily allow access to Jupyter Notebooks from a container by fixing the token and IP address:
@@
where IdentityProvider.token=scoped gives a specific token (this avoids users tracking it in the long logs printed on the terminal), allow-root grants root access for users inside the container volume, ip=0.0.0.0 tells the server to listen on all available network interfaces, and nohup jupyter lab --no-browser & starts Jupyter Lab in the background without attempting to open a browser because launching a browser is not possible on a remote virtual machine that lacks a graphical user interface and is protected by a private IP address. This small code snippet was designed to accelerate research rather than being hung up on infrastructure.